# Robust Multimodal Learning via Cross-Modal Proxy Tokens

### v2

@@ Abstract
Multimodal models often experience a significant performance drop when one or more modalities are missing during inference. To address this challenge, we propose a simple yet effective approach that enhances robustness to missing modalities while maintaining strong performance when all modalities are available. Our method introduces cross-modal proxy tokens (CMPTs), which approximate the class token of a missing modality by attending only to the tokens of the available modality. To efficiently learn the approximation for the missing modality via CMPTs with minimal computational overhead, we employ low-rank adapters in frozen unimodal encoders and jointly optimize an alignment loss with a task-specific loss. Extensive experiments on five multimodal datasets show that our method outperforms state-of-the-art baselines across various missing rates while achieving competitive results in complete-modality settings. Overall, our method offers a flexible and efficient solution for robust multimodal learning. The code and pretrained models will be released on GitHub.
@@ 1 Introduction
Figure 1: Our proposed approach, cross-modal proxy tokens (CMPTs), achieves state-of-the-art performance in both complete and missing modality scenarios. The radar plot shows F1-macro scores on the MM-IMDb dataset, with each axis representing the % availability of image and text modalities. We compare CMPTs against four recent methods [21, 29, 22, 20] using the same experimental setup as  [22, 20].
Figure 2: Illustration of our proposed method. We use pretrained unimodal encoders and adapt them to learn cross-modal proxy tokens (CMPTs) that can approximate missing modality class tokens. We learn the CMPTs and adaptation parameters by jointly optimizing alignment and task-specific objectives. Our approach can accommodate complete and missing modalities during both training and inference and enhances robustness across various missing modality scenarios.
@@
Multimodal learning [35, 52] integrates information from multiple (diverse) input sources to improve the performance on downstream tasks. For example, examining a movie poster image along with a short text synopsis can provide a greater insight into the movie’s type and genre. In recent years, a number of methods have been proposed to effectively combine information from multiple modalities, including early-stage fusion [31], token-level fusion [49], channel exchange [47], and bottleneck mid-fusion [32]. Furthermore, specialized architectures have been proposed for specific multimodal applications, such as vision-language tasks [21, 23, 27], action recognition [51, 56], and segmentation [53, 40]. Most of these approaches assume that all modalities are available for every sample during both training and inference, yielding better performance than unimodal baselines. However, their performance drops sharply when one or more modalities get missing [39, 29].
@@
In practical applications, input modalities can be missing during training and/or inference for various reasons; such as lack of synchronous data capture (limited paired data during training) and sensor failure or privacy concerns (modalities missing during inference). Ma et al. [29] observed that multimodal transformers such as ViLT [21] exhibit a drastic performance drop when one or more modalities are missing during inference as shown in Figure 1 (yellow curve). They developed an optimal fusion strategy to recover the performance loss of multimodal transformers in the missing modality scenarios (purple curve). Subsequently, Lee et al. [22] developed a prompt tuning technique (pink and blue curves), and Kim et al. [20] combined the prompt-based approach with Variance Covariance Invariance regularization (VICReg) (brown curve) to further improve the robustness of multimodal models to missing modalities. Achieving high performance when modalities are missing at varying rates remains a challenge, which we seek to address in this paper.
@@
In this paper, we introduce a simple yet effective framework for multimodal learning that leverages pre-trained unimodal encoders to approximate missing modalities from the available ones. Specifically, we introduce cross-modal proxy tokens (CMPTs) that learn an approximation of the class tokens of the missing modality by attending on all the tokens of the available modality to handle different missing scenarios. As illustrated in Figure 2, our framework integrates pre-trained unimodal encoders into a multimodal learning setting while introducing CMPTs to bridge the performance gap in missing modality scenarios. To ensure efficient learning of CMPTs with minimal additional learnable parameters, we employ low-rank adapters within the frozen unimodal encoders and jointly optimize both an alignment loss and a task-specific loss. The radar plot in Figure 1 demonstrates the superior performance of our approach compared to five existing methods [21, 22, 20, 29] across six different complete and missing modality scenarios on the MM-IMDb dataset, where our CMPT-based model consistently outperforms all baselines. We conduct extensive experiments on five multimodal datasets, comparing our approach against existing state-of-the-art baselines. Results show that our framework achieves state-of-the-art performance in both missing modality (Section 4.3) and complete modality (Section 4.4) scenarios, while maintaining a minimal number of learnable parameters. Ablation studies further confirm that our method significantly improves performance even in severe missing-modality scenarios and provides performance enhancement across most of the classes, leading to an overall performance boost (Section 4.6).
Main contributions of our work are as follows.
@@
  • We propose cross-modal proxy tokens (CMPTs) that learn an approximation of the missing modality from the available modality with minimal adaption of pre-train unimodal encoders via an alignment loss.
@@
  • Our approach outperforms 12 recent baseline methods across four multimodal datasets in various missing-modality scenarios (Section 4.3).
@@ 2 Related Work
Recent works [29, 14, 39] show that performance can degrade significantly when modalities are missing, as multimodal models are not inherently robust to such scenarios. Several approaches have been proposed to address this issue. In this section, we primarily discuss works where the main focus is on making multimodal models robust to missing modalities. More broader related work on multimodal learning is discussed in Supplementary Section S1.
@@
Robust model design is one of the approaches for handling missing modalities. Recently, Shin et al. [41] developed a robust framework utilizing modality masking and knowledge distillation. Wang et al. [44] proposed ShaSpec to impute missing features through modality-specific and modality-shared feature learning. Wang et al. [48] proposed TokenFusion, which dynamically replaces uninformative tokens to enhance robustness. Additionally, different fusion strategies [5, 10, 25] have been developed to improve robustness across various missing-modality scenarios. However, many of these approaches are task-specific, limiting their generalization to other multimodal tasks. In our work, we focus on reusing the pre-trained unimodal encoders, making it flexible for extension to different input modalities and multimodal tasks.
@@
Robust training and model adaptation can enhance missing modality robustness. Recently Reza et al. [39] proposed a parameter-efficient adaptation framework for different multimodal tasks. Nezakati et al. [34] trained a single model to infer missing information by projecting available modalities. Ma et al. [28] introduced SMIL, a Bayesian meta-learning method for flexible handling of missing modalities during both training and testing. Ma et al. [29] proposed a multi-task optimization approach with a differentiable algorithm for optimal fusion strategy selection. Maheshwari et al. [30] developed a semi-supervised framework that leverages unlabeled data to improve robustness, while Wei et al. [50] used a teacher network to transfer multimodal information for improved missing modality robustness. These methods, however, either require extensive adaptation or specialized training procedures. In our approach, we employ lightweight adaptation of the unimodal encoders (using Rank-1 LoRA adapters) and a single extra loss term to explicitly learn cross-modal relationships. Hence, the overall training flow is mostly left unchanged.
@@
Prompt tuning methods improve missing modality performance by incorporating learnable prompts across multiple layers of the model. Lee et al. [22] introduced missing-aware prompts that learns a separate set of prompts for each modality combination. Later, Jang et al. [16] simplified the approach, showing that a single set of prompts per modality can achieve similar performance. More recently, Dai et al. [6] proposed a multi-step adaptive prompt learning method through modality alignment, while Kim and Kim [20] introduced a read-only prompt tuning approach. Although effective, these methods require learning and storing a large number of prompts, typically 16 prompts [22, 16] per modality per layer, for each missing modality scenario. In contrast, our method uses a single cross-modal proxy token per modality to estimate the missing modality feature from the available modality, thereby reducing the complexity. We carry out extensive experimental comparisons with all the above prompt-based techniques to demonstrate superior performance across various missing modality scenarios. Furthermore, our approach is efficient, uses few learnable parameters, and is robust to missing modalities in both training and testing which we discuss in Section 4.
@@ 3 Proposed Approach
In this paper, we propose a new framework to train multimodal systems such that they retain their performance under different scenarios of missing modalities during inference. Furthermore, we seek to develop a simple and efficient framework that can be used for different multimodal tasks and datasets without any specialized (task-specific) changes. To achieve these objectives, we propose to leverage modality-specific (transformer-based) encoders that are pre-trained on large datasets and align the features across modalities using proxy tokens, as illustrated in Figure 2.
@@ 3.1 Multimodal embedding tokens
Let us consider a multimodal dataset containing two modalities =$\mathcal{M}=\{m_{1},m_{2}\}$. We denote the available dataset as $\mathcal{D}=\{\mathcal{D}_{c},\mathcal{D}_{m_{1}},\mathcal{D}_{m_{2}}\}$, which can be further divided into three subsets: $\mathcal{D}_{c}=\{x_{m_{1}},x_{m_{2}},y\}$ where both modalities are present with corresponding output labels $y$; $\mathcal{D}_{m_{1}}=\{x_{m_{1}},\varnothing,y\}$ where only modality $m_{1}$ $(x_{m_{1}})$ is available with corresponding labels $y$; $\mathcal{D}_{m_{2}}=\{\varnothing,x_{m_{2}},y\}$ where only modality $m_{2}$ $(x_{m_{2}})$ is available with corresponding labels $y$. Here, $\varnothing$ denotes a placeholder for the missing modality (e.g., empty string for text, zeros for image, audio, or video) following [22, 39].
For each input modality, we assume that a transformer-based encoder, pre-trained on some large dataset, is available. Specifically, each modality $m\in\mathcal{M}$ is processed through a pre-trained embedding layer, $\text{EmbeddingLayer}_{m}$, with parameters $\Theta_{e,m}$. The input $x_{m}$ for each modality is passed through its corresponding embedding layer to generate modality-specific tokens as
@@
3.2 Cross-modal proxy tokens
@@
Note that the proposed CMPT serves as an extra token, similar to the class token [8, 7], but with a slightly orthogonal purpose. Class token attends to all the tokens of a given modality to aggregate information and learn a rich feature representation of that modality. In contrast, CMPT attends to all the tokens of a given modality to learn an approximation of the class token of the other modality (e.g., $\mathcal{T}_{\text{CMPT},m_{1}}$ approximates $\mathcal{T}_{\text{CLS},m_{2}}$ while attending to $\mathcal{T}_{m_{1}}$).
@@ 3.3 Learn CMPTs by alignment and adaptation
For each modality, we assume that a pretrained transformer encoder $f_{m}$ with parameters $\Theta_{f,m}$ is available. The encoder for each modality needs to be updated so that they can learn the corresponding CMPTs. To keep the number of trainable parameters minimal, we use low-rank adaptation (LoRA) [15] to learn a small number of parameters $\Delta_{f,m}$. The modified embedding tokens for modality $m$, $\tilde{\mathcal{T}}_{m}$, are then passed through the adapted encoder for modality $m$ to get the final output tokens as
@@
where $\hat{\mathcal{T}}_{m}\in\mathbb{R}^{(N+2)\times d}$ represents the CMPT, CLS, and remaining tokens from the last encoder layer. Note that the pretrained encoder parameters $\Theta_{f,m}$ remain frozen while we learn $\Delta_{f,m}$.
@@ 3.4 Gating module and fusion
Finally, we learn the low-rank factors for each encoder ($(\{\Delta_{f,m}\})$ and classifier head $(\Theta_{h})$ by minimizing the task-specific loss $\mathcal{L}_{\text{task}}$ combined with the alignment loss:
@@
where $\lambda$ is a hyperparameter that we set to $\lambda=0.20$ in all our experiments.
Our model can handle missing modalities during both training and inference. During training on complete modality data, we apply random modality dropout, where we either use both modalities or randomly drop one in each iteration. This way, the model learns to make correct predictions in missing modality scenarios using the CMPTs from the available modality. When training data contains missing modalities, we learn the CMPTs by optimizing the alignment loss using only the samples with complete modalities, as shown in (4). For missing modalities, we employ placeholders (e.g., an empty string for text and zeros for images, audio, and video) following standard practices [22, 39], and use the CMPT from the available modality as an approximation to make predictions, as described in (5). To enhance training efficiency, we leverage LoRA [15] with a rank of 1 to fine-tune the encoders in a parameter-efficient manner. LoRA layers are inserted after the query, key, value, and output layers in every attention blocks within the encoders. Our experiments demonstrate that LoRA with rank 1 is sufficient to achieve better performance compared to existing state-of-the-art methods, as discussed in Section 4.

### v4

@@ Abstract
Multimodal models often experience a significant performance drop when one or more modalities are missing during inference. To address this challenge, we propose a simple yet effective approach that enhances robustness to missing modalities while maintaining strong performance when all modalities are available. Our method introduces cross-modal proxy tokens (CMPTs), which approximate the class token of a missing modality by attending only to the tokens of the available modality without requiring explicit modality generation or auxiliary networks. To efficiently learn these approximations with minimal computational overhead, we employ low-rank adapters in frozen unimodal encoders and jointly optimize an alignment loss with a task-specific loss. Extensive experiments on five multimodal datasets show that our method outperforms state-of-the-art baselines across various missing rates while achieving competitive results in complete-modality settings. Overall, our method offers a flexible and efficient solution for robust multimodal learning. The code for this paper is available at: https://github.com/CSIPlab/Cross-Modal-Proxy-Tokens.
@@ 1 Introduction
Multimodal learning (Ngiam et al., 2011; Xu et al., 2023) integrates information from multiple (diverse) input sources to improve the performance on downstream tasks. For example, examining a movie poster image along with a short text synopsis can provide a greater insight into the movie’s type and genre. In recent years, a number of methods have been proposed to effectively combine information from multiple modalities, including early-stage fusion (Mo & Morgado, 2024), token-level fusion (Wang et al., 2022b), channel exchange (Wang et al., 2020b), and bottleneck mid-fusion (Nagrani et al., 2021a). Furthermore, specialized architectures have been proposed for specific multimodal applications, such as vision-language tasks (Kim et al., 2021; Li et al., 2019; Lu et al., 2019), action recognition (Woo et al., 2023; Zhou et al., 2022), and segmentation (Zhang et al., 2023a; Reza et al., 2024b). Most of these approaches assume that all modalities are available for every sample during both training and inference, yielding better performance than unimodal baselines. However, their performance drops sharply when one or more modalities get missing (Reza et al., 2024a; Ma et al., 2022).
@@
In practical applications, input modalities can be missing during training and/or inference for various reasons; such as lack of synchronous data capture (limited paired data during training) and sensor failure or privacy concerns (modalities missing during inference). Ma et al. (2022) observed that multimodal transformers such as ViLT (Kim et al., 2021) exhibit a drastic performance drop when one or more modalities are missing during inference as shown in Figure 1(b) (yellow curve). They developed an optimal fusion strategy to recover the performance loss of multimodal transformers in the missing modality scenarios (purple curve). Subsequently, Lee et al. (2023) developed a prompt tuning technique (pink and blue curves), and Kim & Kim (2024) combined the prompt-based approach with Variance Covariance Invariance regularization (VICReg) (brown curve) to further improve the robustness of multimodal models to missing modalities. Despite these efforts, achieving consistently strong performance under varying degrees of missing modalities remains an open challenge. In this work, we aim to improve robustness across a wide range of missing modality conditions while maintaining strong performance when all modalities are available.
(a) Illustration of our proposed method.
(b) Performance comparison.
Figure 1: (a) We introduce Cross-Modal Proxy Tokens (CMPTs), a novel approach to address missing modality challenges. CMPTs effectively learn to approximate missing modality class tokens by adapting pretrained encoders through a joint optimization of alignment and task-specific objectives. Our approach accommodates both complete and missing modalities during training and inference, thereby enhancing robustness across varying missing modality scenarios. (b) CMPTs achieve state-of-the-art performance, consistently outperforming recent baseline methods in both complete and missing modality scenarios. Following the experimental setup of Lee et al. (2023); Kim & Kim (2024), the radar plot illustrates F1-macro scores on the MM-IMDb dataset across varying modality availability.
@@
In this paper, we introduce a simple yet effective framework to approximate missing modalities from the available ones for robust multimodal learning. Specifically, we introduce cross-modal proxy tokens (CMPTs), which learn to approximate the class tokens of missing modalities by attending to all tokens of the available modality, enabling effective handling of various missing modality scenarios. As illustrated in Figure 1(a), our framework integrates pre-trained unimodal encoders into a multimodal learning setting while introducing CMPTs to partially bridge the performance gap in missing modality scenarios. To ensure efficient learning of CMPTs with minimal additional learnable parameters, we employ low-rank adapters within the frozen unimodal encoders and jointly optimize both an alignment loss and a task-specific loss. The radar plot in Figure 1(b) demonstrates the superior performance of our approach compared to five existing methods (Kim et al., 2021; Lee et al., 2023; Kim & Kim, 2024; Ma et al., 2022) across six different complete and missing modality scenarios on the MM-IMDb dataset. Our CMPT-based model consistently outperforms all baselines across all the scenarios. We conduct extensive experiments on five multimodal datasets, comparing our approach against existing state-of-the-art baselines. Results show that our framework achieves state-of-the-art performance in both missing modality (Section 4.3) and complete modality (Section 4.4) scenarios, while maintaining a minimal number of learnable parameters. Ablation studies further confirm that our method significantly improves performance even in severe missing-modality scenarios and provides performance enhancement across most of the classes, leading to an overall performance boost (Section 4.6). Main contributions of our work are as follows.
@@
  • We propose CMPTs to effectively learn an approximation of the missing modality from the available modality with minimal adaption of pre-trained unimodal encoders via an alignment loss.
@@
  • Our approach outperforms 12 recent baseline methods across four multimodal datasets in various missing-modality scenarios, demonstrating state-of-the-art performance (Section 4.3).
@@ 2 Related Work
Recent works (Ma et al., 2022; Hazarika et al., 2022; Reza et al., 2024a) show that performance can degrade significantly when modalities are missing, as multimodal models are not inherently robust to such scenarios. Several approaches have been proposed to address this issue. In this section, we primarily discuss works where the main focus is on making multimodal models robust to missing modalities. More broader related work on multimodal learning is discussed in appendix Section A1.
@@
Robust model design is one of the approaches for handling missing modalities. Recently, Shin et al. (2024) developed a robust framework utilizing modality masking and knowledge distillation. Wang et al. (2023) proposed ShaSpec to impute missing features through modality-specific and modality-shared feature learning. Wang et al. (2022a) proposed TokenFusion, which dynamically replaces uninformative tokens to enhance robustness. Additionally, different fusion strategies (Choi & Lee, 2019; Fan et al., 2023a; Lin et al., 2023) have been developed to improve robustness across various missing-modality scenarios. However, many of these approaches are task-specific, limiting their generalization to other multimodal tasks. In our work, we focus on reusing pre-trained unimodal encoders, making it flexible for extension to different input modalities and multimodal tasks.
@@
Robust training and model adaptation can enhance missing modality robustness. Recently, Reza et al. (2024a) proposed a parameter-efficient adaptation framework for different multimodal tasks. Shi et al. (2024) introduced a modality-aware low-rank adaptation method (MoRA) that uses a shared down-projection layer to map inputs to a low-dimensional space, and then applies modality-specific up-projections for missing modality adaptation. Nezakati et al. (2024) trained a single model to infer missing information by projecting available modalities. Ma et al. (2021) introduced SMIL, a Bayesian meta-learning method for flexible handling of missing modalities during both training and testing. Ma et al. (2022) proposed a multi-task optimization approach with a differentiable algorithm for optimal fusion strategy selection. Maheshwari et al. (2024) developed a semi-supervised framework that leverages unlabeled data to improve robustness, while Wei et al. (2023) used a teacher network to transfer multimodal information for improved missing modality robustness. These methods, however, either require extensive adaptation or specialized training procedures. In our approach, we employ lightweight adaptation of unimodal encoders (using Rank-1 LoRA adapters) and a single extra loss term to explicitly learn cross-modal relationships. Hence, the overall training flow is mostly left unchanged.
@@
Prompt tuning methods improve missing modality performance by incorporating learnable prompts across multiple layers of the model. Lee et al. (2023) introduced missing-aware prompts that learns a separate set of prompts for each modality combination. Later, Jang et al. (2024) simplified the approach, showing that a single set of prompts per modality can achieve similar performance. More recently, Dai et al. (2024) proposed a multi-step adaptive prompt learning method through modality alignment, while Kim & Kim (2024) introduced a read-only prompt tuning approach with Variance Covariance Invariance regularization. Although effective, these methods require learning and storing a large number of prompts, typically 16 prompts (Lee et al., 2023; Jang et al., 2024) per modality per layer, for each missing modality scenario. In contrast, our method uses a single cross-modal proxy token per modality to estimate the missing modality feature from the available modality, thereby reducing the complexity.
We carry out extensive experimental comparisons with all the above prompt-based techniques to demonstrate superior performance across various missing modality scenarios. Furthermore, our approach is efficient, uses few learnable parameters, and is robust to missing modalities in both training and testing which we discuss in Section 4.
@@ 3 Proposed Approach
In this paper, we propose a new framework to train multimodal systems such that they retain strong performance under different scenarios of missing modalities during inference. Furthermore, we seek to develop a simple and efficient framework that can be used for different multimodal tasks and datasets without any specialized (task-specific) changes. To achieve these objectives, we propose cross-modal proxy tokens (CMPTs), which effectively learn to approximate the class tokens of missing modalities by adapting pretrained, modality-specific (transformer-based) encoders through a joint optimization of alignment and task-specific objectives, as illustrated in Figure 1(a).
@@ 3.1 Multimodal embedding tokens
Let us consider a multimodal dataset containing two modalities $\mathcal{M}=\{m_{1},m_{2}\}$. We denote the available dataset as $\mathcal{D}=\{\mathcal{D}_{c},\mathcal{D}_{m_{1}},\mathcal{D}_{m_{2}}\}$, which can be further divided into three subsets: $\mathcal{D}_{c}=\{x_{m_{1}},x_{m_{2}},y\}$ where both modalities are present with corresponding output labels $y$; $\mathcal{D}_{m_{1}}=\{x_{m_{1}},\varnothing,y\}$ where only modality $m_{1}$ $(x_{m_{1}})$ is available with corresponding labels $y$; $\mathcal{D}_{m_{2}}=\{\varnothing,x_{m_{2}},y\}$ where only modality $m_{2}$ $(x_{m_{2}})$ is available with corresponding labels $y$. Here, $\varnothing$ denotes a placeholder for the missing modality (e.g., empty string for text, zeros for image, audio, or video) following prior works (Lee et al., 2023; Reza et al., 2024a).
For each input modality, we assume that a transformer-based encoder, pre-trained on some large dataset, is available. This is a common and effective practice in modern multimodal learning, particularly given the wide variety of high-performing, pre-trained models now readily accessible (Devlin et al., 2019; Dosovitskiy et al., 2021; Gong et al., 2021). Leveraging these modality-specific encoders trained on extensive data leads to improved performance and faster convergence (Han et al., 2023; Xu et al., 2023). Specifically, each modality $m\in\mathcal{M}$ is processed through a pre-trained embedding layer, $\text{EmbeddingLayer}_{m}$, with parameters $\Theta_{e,m}$. The input $x_{m}$ for each modality is passed through its corresponding embedding layer to generate modality-specific tokens as
@@
3.2 Cross-Modal Proxy Tokens (CMPTs)
@@
Note that the proposed CMPT serves as an extra token, similar to the class token (Dosovitskiy et al., 2021; Devlin et al., 2019; Gong et al., 2021), but with a complementary purpose. Class token attends to all the tokens of a given modality to aggregate information and learn a rich feature representation of that modality. In contrast, CMPT attends to all the tokens of a given modality to learn an approximation of the class token of the other modality (e.g., $\mathcal{T}_{\text{CMPT},m_{1}}$ approximates $\mathcal{T}_{\text{CLS},m_{2}}$ while attending to $\mathcal{T}_{m_{1}}$).
@@ 3.3 Learn CMPTs by alignment and adaptation
For each modality, we assume that a pretrained transformer encoder $f_{m}$ with parameters $\Theta_{f,m}$ is available. The encoder for each modality needs to be updated so that they can learn the corresponding CMPTs. To keep the number of trainable parameters minimal, we use low-rank adaptation (LoRA) (Hu et al., 2022) to learn a small number of parameters $\Delta_{f,m}$. The modified embedding tokens for modality $m$, $\tilde{\mathcal{T}}_{m}$, are then passed through the adapted encoder for modality $m$ to get the final output tokens as
@@
where $\hat{\mathcal{T}}_{m}\in\mathbb{R}^{(N+2)\times d}$ represents the CMPT, CLS, and remaining tokens from the last encoder layer. Note that the pretrained encoder parameters $\Theta_{f,m}$ remain frozen while we only learn $\Delta_{f,m}$.
@@ 3.4 Gating module and fusion
Finally, we learn the low-rank factors for each encoder $(\{\Delta_{f,m}\})$ and classifier head $(\Theta_{h})$ by minimizing the task-specific loss $\mathcal{L}_{\text{task}}$, which for our purposes is a standard cross-entropy loss for classification tasks, combined with the alignment loss:
@@
where $\lambda$ is a hyperparameter that we set to $\lambda=0.20$ in all our experiments. As discussed in Section 4.6.3, this value provides a balance in performance across different scenarios.
3.5 Handling missing modalities
Our model can handle missing modalities during both training and inference. During training on complete modality data, we apply random modality dropout, where we either use both modalities or randomly drop one in each iteration. This way, the model learns to make correct predictions in missing modality scenarios using the CMPTs from the available modality. When training data contains missing modalities, we learn the CMPTs by optimizing the alignment loss using only the samples with complete modalities, as shown in equation 4. For missing modalities, we employ placeholders (e.g., an empty string for text and zeros for images, audio, and video) following standard practices (Lee et al., 2023; Reza et al., 2024a), and use the CMPT from the available modality as an approximation to make predictions, as described in equation 5.
To enhance training efficiency, we leverage LoRA (Hu et al., 2022) with a rank of 1 to fine-tune the encoders in a parameter-efficient manner. LoRA layers are inserted after the query, key, value, and output layers in every attention blocks within the encoders. Our experiments demonstrate that LoRA with rank 1 is sufficient to achieve better performance compared to existing state-of-the-art methods, as discussed in Section 4.4 and Section A4. Our method is also compatible with newer parameter-efficient adaptation techniques such as DoRA (Liu et al., 2024) and VeRA (Kopiczko et al., 2024), as discussed in Section A10.